OpenStudy (anonymous):

A bag contains 3 yellow tennis balls and 4 green tennis balls. Gordy will select one tennis ball, put it in his pocket, and then select a second tennis ball. Which expression gives the probability that he will select two green tennis balls? A.4/7*3/6 B.4/7*4/6 C.4/7*3/7 D.4/7*4/7

OpenStudy (anonymous):

c prob

OpenStudy (agl202):

There are 3 + 4 = 7 tennis balls to start with and 4 of them are green. Therefore the probability of the first being green = 4/7. Once he puts this ball in his ocket there are only 6 balls left to select from and now only three are green. Remember the other green ball is in his pocket. Therefore the probability that the second is green also = 3/6. Finally as you are asked the probability that BOTH are green, multiply the individual probabilities together 4/7 * 3/6 = Answer A Hopefully u understand how I'm getting this all... :D